# A Hierarchical Framework for Measuring Scientific Paper Innovation via Large Language Models

### v1

@@ III-B HSPIM Framework and Workflow
III-B1 Step-1: Section Classification
@@
III-B2 Step-2: Question-Answering Augmentation
@@
[…]
@@
III-B3 Step-3: Innovation Scoring
@@
The formula for calculating the predicted innovation score $\hat{\textit{Innovation}}$ is given by $\hat{\textit{Innovation}}_{i}=\sum_{k=1}^{l_{i}}\textit{Confidence}_{ik}\cdot\textit{Novelty}_{ik}/\sum_{k=1}^{l_{i}}\textit{Confidence}_{ik}$, where $i$ denotes the $i$-th paper in the dataset, and $l_{i}$ represents the total number of text chunks of the $i$-th paper.
III-B4 Step-4: Evaluate Predicted Innovation
We compute the difference between the HSPIM output innovation scores with the ground-truth peer review scores in the dataset. Our objective is to minimize the RMSE of innovation scores. Since the current peer reviews do not provide a direct innovation score, we instead focus on two aspect scoring: originality and soundness/correctness, both rated on a scale of {1, 2, 3, 4, 5}. Originality reflects the novelty, and soundness/correctness scores the rigor, validity, and contribution to the field. We define innovation as a combination of originality and soundness/correctness. A paper with a high originality score but a low soundness/correctness score may have methodological flaws or unreliable conclusions, reducing its overall innovation. Thus, the peer review innovation score $\textit{Innovation}_{i}$ for paper $p_{i}$ is computed as the average of originality and soundness scores across all its reviews:
@@ III-B4 Step-4: Evaluate Predicted Innovation
III-B5 Step-5: Optimize Question Prompt
@@
In our GA implementation, we define an individual as a question-prompt combination. As shown in Fig. 2, each individual consists of one common question and $s$ specific questions. All questions are selected from their corresponding question sets. We initialize the population by randomly generating $p$ individuals for each paper. To compare the effectiveness of different question-prompt combinations (individuals), each paper is evaluated using $p$ different question combinations, producing $p$ final innovation scores. The GA prompt optimizer calculates the fitness of all $p$ individuals using RMSE and ranks them accordingly. The top-performing individuals are selected as parents for the next generation while keeping elite individuals.
@@
III-C Framework Comparison
As described in Section III-B, we progressively construct the hierarchical scientific paper innovation measurement (HSPIM) framework, which consists of Step-1 to Step-5.
@@ III-C Framework Comparison
III-D Multi-prompt Optimization Strategies in HSPIM
@@
In multi-prompt optimization, question prompts evolve similarly to genetic algorithms, where specific questions function analogously to structural genes affecting distinct traits, and common questions act as regulatory genes influencing broader characteristics [44, 45]. Each specific question targets precisely one section type, akin to a mutation in a structural gene causing targeted phenotypic changes. In contrast, the common question governs global aspects of innovation evaluation, mirroring the role of regulatory genes in coordinating the expression of multiple genes simultaneously [44].
@@
IV Theoretical Analysis of HSPIM
In this section, we provide a theoretical analysis of our HSPIM framework, focusing on: (1) the multi-level weighted scoring of section-based evaluations, and (2) the convergence properties of the genetic algorithm-based prompt optimization. Our notation and definitions follow Section III-A, but we make slight modifications to accommodate the revised bounded interval and unbiasedness assumptions for both $\textit{Novelty}_{ik}$ and $\textit{Confidence}_{ik}$.
IV-A Preliminaries and Assumptions
@@
    Section-Based Unbiasedness (in expectation): For each chunk $t_{ik}$, the random variables $\textit{Novelty}_{ik}$ and $\textit{Confidence}_{ik}$ satisfy $\mathbb{E}[\textit{Novelty}_{ik}]=n_{ik}^{*},\;\mathbb{E}[\textit{Confidence}_{ik}]=c_{ik}^{*},$ and $\textit{Novelty}_{ik}$ is independent of $\textit{Confidence}_{ik}$. Here, $n_{ik}^{*}$ and $c_{ik}^{*}$ represent “true” local novelty and confidence values, respectively.
@@
IV-B Convergence of the Weighted Scoring Mechanism
@@
Theorem IV.1 (Unbiasedness of the weighted scoring function).
@@
Since $\textit{Novelty}_{ik}$ and $\textit{Confidence}_{ik}$ are independent, we have $\mathbb{E}[\textit{Confidence}_{ik}\,\textit{Novelty}_{ik}]=\mathbb{E}[\textit{Confidence}_{ik}]\,\mathbb{E}[\textit{Novelty}_{ik}]=c_{ik}^{*}\,n_{ik}^{*}$. Substituting completes the proof. ∎
Discussion. If local chunk-level novelty and confidence are unbiased in expectation, the final weighted average preserves unbiasedness for overall innovation. Each chunk focuses on a narrower aspect of novelty, weighted by its confidence.
IV-C Convergence of GA-Based Prompt Optimization
@@
Definition IV.2 (Discrete Prompt Combination).
@@
Theorem IV.3 (GA Convergence on Finite Space).
@@
IV-D Overall Effectiveness and Convergence Discussion
Combining Theorem IV.1 and Theorem IV.3 shows:
@@ IV-D Overall Effectiveness and Convergence Discussion
V Experiment
@@
TABLE II: Statistics of peer review datasets.
@@
V-A Datasets
We choose scientific peer review datasets from ACL-2017, CoNLL-2016, and ICLR-2017 in PeerRead [17], and the COLING-2020 dataset from NLPeer [18]. Table II shows that ICLR-2017 belongs to the ML/AI domain, while the others are from the NLP/CL domain. As describe in Step-4 of Section III-B, these datasets provide numeric aspect scores of originality and soundness/correctness to form the ground-truth innovation score. Note that originality scores are mainly available in NLP-related conferences. For ICLR-2017, originality scores were manually annotated by the PeerRead authors and validated through the consistency re-annotation experiment. All labeled scores range from 1 to 5 (integers only).
V-B Metrics
@@
V-C Baselines
V-D Hyperparameters and Experimental Setting
[…]
Supervised Baseline Settings: We use the PeerRead setup for deep learning models and take the first 1000 tokens of each paper as input. We select only the originality and soundness/correctness aspect scores for training. The training epochs are set to 10 for CNN, LSTM, and Transformer. For CNN and LSTM, we set the number of layers to 2, the dropout rate to 0.5, the learning rate to 0.001, and the batch size to 64. For the Transformer encoder, we use 12 attention heads, a hidden size of 768, a learning rate of 2e-5, and a batch size of 8. All results are averaged over three runs.
TABLE III: Performance comparison among zero-shot HSPIM, supervised deep learning models, and other zero-shot LLMs
@@ V-D Hyperparameters and Experimental Setting
The bold values represent the best results.
V-E Main Results
This section presents the main experimental results. We compare our HSPIM framework with baseline models on four datasets using RMSE and MAE as evaluation metrics. In the following results, HSPIMtwo-step and HSPIMjoint denote the two-step and joint optimization for HSPIM, respectively. HSPIMpruning=3 selects three section types ($s^{\prime}=3$) to compute the innovation scores, following Section III-D. We use HSPIMjoint as the default framework to represent HSPIM with prompt optimization in the following experiments.
V-E1 Zero-shot HSPIM vs. Supervised DL Models and Zero-shot LLMs
Table III shows that HSPIM framework outperforms both supervised baseline models and zero-shot baseline LLMs in terms of RMSE and MAE. Compared to one-step zero-shot prompting on full papers, HSPIM based on paper-to-sections-to-QAs decomposition reduces the average RMSE by 35.4% (from 0.916 to 0.592) for DeepSeek and by 17.3% (from 0.770 to 0.634) for GPT. Besides, our HSPIM outperforms all supervised baselines. This proves that the zero-shot approach can generalize well. For instance, DeepSeek-HSPIMpruning=3 achieves a 6.2% performance improvement over the supervised Transformer.
Additionally, we observe the following: (i) DeepSeek-HSPIM achieves better overall performance than GPT-HSPIM, which is reasonable because DeepSeek-V3 has stronger fundamental capabilities. (ii) HSPIMpruning=3 achieves the best results among the three optimization strategies. Fig. 5 shows the top-3 selected section types across different datasets and LLMs. We observe that Abstract, Analysis Theory, Experiment Analysis, and Conclusion frequently appear. It indicates these sections contain important novelty explanations and method contributions for measuring innovation. We believe section pruning helps to filter out potential noisy section types. In addition, it guides human reviewers to focus on specific sections that most significantly contribute to innovation analysis. (iii) HSPIM achieves relatively lower performance on the ACL-2017 dataset due to dataset differences. ACL-2017 has a higher average ground-truth innovation score (4.321) on the test set than ICLR-2017 (3.696), CoNLL-2016 (3.361), and COLING-2020 (3.615). Supervised baselines perform better since both training and test sets in ACL-2017 have high scores. LLM baselines (DeepSeek-V3 and GPT-4o mini) also perform well on ACL-2017 because LLMs tend to assign higher scores to longer texts [38]. This tendency matches the scoring distribution of ACL-2017. However, LLM baselines perform poorly on other datasets.
V-E2 Zero-Shot HSPIM vs. DL Models (Transfer from COLING-2020)
Table IV compares HSPIM with DL models trained on COLING-2020 and evaluated on the remaining three datasets. This setting provides a fairer comparison. HSPIM significantly outperforms DL models on ICLR-2017, CoNLL-2016, and ACL-2017. Compared to the best results of CNN, LSTM, and Transformer, HSPIM achieves average improvements of 44.0%, 41.8%, and 33.1%, respectively. These results confirm the model generalization ability of the zero-shot HSPIM framework.
V-E3 HSPIM Framework Comparison
Table V shows the performance comparison of three framework variants: section-based scientific paper innovation measurement (SSPIM), HSPIM using the initial prompt combination (naive HSPIM), and HSPIM with GA-based joint optimization strategy (HSPIMjoint). We report results on both the training set and the test set. We observe a consistent decrease in RMSE from SSPIM to naive HSPIM, and further to HSPIMjoint, which highlights the importance of each component. Specifically, naive HSPIM outperforms SSPIM, which shows the effectiveness of QA augmentation. HSPIMjoint improves naive HSPIM, which confirms the benefit of multi-prompt optimization.
TABLE IV: Performance comparison on ICLR-2017, CoNLL-2016, and ACL-2017: zero-shot HSPIM vs. DL models trained on COLING-2020
@@ V-E3 HSPIM Framework Comparison
The bold values represent the best results.
TABLE V: RMSE comparison across our proposed frameworks

### v2

@@ 3.2 HSPIM Framework and Workflow
3.2.1 Step-1: Section Classification
@@
3.2.2 Step-2: Question-Answering Augmentation
@@
$\textit{section question}=\textit{specific question}+\textit{common question}.$ (6)
@@
3.2.3 Step-3: Innovation Scoring
@@
The formula for calculating the predicted innovation score $\hat{\textit{Innovation}}$ is given by $\hat{\textit{Innovation}}_{i}=\sum_{k=1}^{l_{i}}\textit{Confidence}_{ik}\cdot\textit{Novelty}_{ik}/\sum_{k=1}^{l_{i}}\textit{Confidence}_{ik}$, where $i$ denotes the $i$-th paper in the dataset, and $l_{i}$ represents the total number of text chunks of the $i$-th paper. Based on innovation theory, we also introduce the innovation attributes of contribution and feasibility, and define HSPIM+ in Definition 3.2.
3.2.4 Step-4: Evaluate Predicted Innovation
We compute the difference between the HSPIM output innovation scores with the ground-truth peer review scores in the dataset. Our objective is to minimize the RMSE of innovation scores. Since the current peer reviews do not provide a direct innovation score, we instead focus on two attribute scoring: originality and soundness/correctness, both rated on a scale of {1, 2, 3, 4, 5}. Originality reflects the novelty, and we use soundness/correctness as a proxy for practical value. We define innovation as a combination of originality and soundness/correctness. A paper with a high originality score but a low soundness/correctness score may have methodological flaws or unreliable conclusions, reducing its overall innovation. Thus, the peer review innovation score $\textit{Innovation}_{i}$ for paper $p_{i}$ is computed as the average of originality and soundness scores across all its reviews:
@@ 3.2.4 Step-4: Evaluate Predicted Innovation
3.2.5 Step-5: Optimize Question Prompt
@@
In our GA implementation, we define an individual as a question-prompt combination. As shown in Fig. 3, each individual consists of one common question and $s$ specific questions. All questions are selected from their corresponding question sets. We initialize the population by randomly generating $p$ individuals for each paper. To compare the effectiveness of different question-prompt combinations (individuals), each paper is evaluated using $p$ different question combinations, producing $p$ final innovation scores. The GA prompt optimizer calculates the fitness of all $p$ individuals using RMSE and ranks them accordingly. Specifically, the fitness function is defined as the RMSE between the predicted and ground-truth innovation scores on the validation set. The top-performing individuals are selected as parents for the next generation while keeping elite individuals.
@@
3.3 Framework Comparison
As described in Section 3.2, we progressively construct the hierarchical scientific paper innovation measurement (HSPIM) framework, which consists of Step-1 to Step-5.
@@ 3.3 Framework Comparison
3.4 Multi-prompt Optimization Strategies in HSPIM
@@
In multi-prompt optimization, question prompts evolve similarly to genetic algorithms, where specific questions function analogously to structural genes affecting distinct traits, and common questions act as regulatory genes influencing broader characteristics [37, 38]. Each specific question targets precisely one section type, akin to a mutation in a structural gene causing targeted phenotypic changes. In contrast, the common question governs global attributes of innovation evaluation, mirroring the role of regulatory genes in coordinating the expression of multiple genes simultaneously [37].
@@
4 Theoretical Analysis of HSPIM
In this section, we provide a theoretical analysis of our HSPIM framework, focusing on the unbiasedness of the weighted scoring function and the convergence properties of the genetic algorithm-based prompt optimization. Our notation and definitions follow Section 3.1, but we make slight modifications to accommodate the revised bounded interval and unbiasedness assumptions for both innovation attributes and $\textit{Confidence}_{ik}$.
4.1 Preliminaries and Assumptions
@@
    Section-Based Unbiasedness (in expectation): For each chunk $t_{ik}$, the random variables $\textit{Novelty}_{ik}$ and $\textit{Confidence}_{ik}$ satisfy $\mathbb{E}[\textit{Novelty}_{ik}]\allowbreak=n_{ik}^{*},\;\mathbb{E}[\textit{Confidence}_{ik}]=c_{ik}^{*}.$ $\textit{Novelty}_{ik}$ is independent of $\textit{Confidence}_{ik}$. Here, $n_{ik}^{*}$ and $c_{ik}^{*}$ represent “true” local novelty and confidence values, respectively.
@@
4.2 Convergence of the Weighted Scoring Mechanism
@@
Theorem 4.1 (Unbiasedness of the weighted scoring function).
@@
Since $\textit{Novelty}_{ik}$ and $\textit{Confidence}_{ik}$ are independent, we have:
$\mathbb{E}[\textit{Confidence}_{ik}\,\textit{Novelty}_{ik}]$ $=\mathbb{E}[\textit{Confidence}_{ik}]\,\mathbb{E}[\textit{Novelty}_{ik}]$ (11)
$=c_{ik}^{*}\,n_{ik}^{*}.$
Substituting completes the proof. ∎
Discussion. If local chunk-level novelty and confidence are unbiased in expectation, the final weighted average preserves unbiasedness for overall innovation. Each chunk focuses on a narrower attribute of novelty, weighted by its confidence.
4.3 Convergence of GA-Based Prompt Optimization
@@
Definition 4.2 (Discrete Prompt Combination).
@@
Theorem 4.3 (GA Convergence on Finite Space).
@@
4.4 Overall Effectiveness and Convergence Discussion
Combining Theorem 4.1 and Theorem 4.3 shows:
  • 1.
@@ 4.4 Overall Effectiveness and Convergence Discussion
4.5 Theoretical Justification of HSPIM+
To capture the composite nature of innovation, HSPIM+ models it as a function of three core attributes: novelty, contribution, and feasibility. The mathematical structure is:
$\textit{Innovation}=f(\textit{Novelty},\,\textit{Contribution},\,\textit{Feasibility}),$ (15)
where $f$ is the mapping defined by a $p$-norm with linear normalization (see Definition 3.2).
We make the following assumptions for all papers $p_{i}$, chunks $t_{ik}$, and innovation attributes:
  • (B1)
    All innovation attribute scores (novelty, contribution, feasibility) and their confidence scores are bounded and strictly positive.
  • (B2)
    Each attribute score is unbiased in expectation and independent of its corresponding confidence score.
  • (B3)
    Innovation attribute scores and confidence scores are independent across section chunks.
  • (B4)
    Idealized Independence: For each chunk, the innovation attribute scores and their confidence scores are mutually independent across novelty, contribution, and feasibility; that is, there is no dependence among the three attributes in either scoring or confidence assignment.
Remark. Assumption (B4) idealizes the three attributes as fully independent for analytical tractability. In practice, some correlation may exist (e.g., high novelty may reduce feasibility), but we follow standard practice in multi-attribute aggregation.
Theorem 4.4 (Unbiasedness of Norm-Based HSPIM+).
Let $\hat{\textit{Innovation}}_{i}^{(p)}$ be the innovation score for paper $p_{i}$ computed by HSPIM+ using the $p$-norm aggregation and linear normalization, as defined in Definition 3.2, under assumptions (B1)–(B4).
Then, the predicted vector $\mathbf{v}_{i}=[\textit{Novelty}_{i},\,\textit{Contribution}_{i},\,\textit{Feasibility}_{i}]$ satisfies:
$\mathbb{E}[\textit{Novelty}_{i}]$ $=n_{i}^{*},$ (16)
$\mathbb{E}[\textit{Contribution}_{i}]$ $=d_{i}^{*},$ (17)
$\mathbb{E}[\textit{Feasibility}_{i}]$ $=f_{i}^{*},$ (18)
where $n_{i}^{*}$, $d_{i}^{*}$, $f_{i}^{*}$ are the true paper-level values for the three attributes, computed as the confidence-weighted means of the corresponding true section-level values.
Moreover, the overall innovation score $\hat{\textit{Innovation}}_{i}^{(p)}$ is an unbiased, monotonic transformation of these means via the $p$-norm and normalization:
$\mathbb{E}[\hat{\textit{Innovation}}_{i}^{(p)}]=1+4\cdot\frac{\|[n_{i}^{*},d_{i}^{*},f_{i}^{*}]\|_{p}-x_{\min}}{x_{\max}-x_{\min}},$ (19)
where $x_{\min}$ and $x_{\max}$ are the theoretical minimum and maximum values of the $p$-norm in $1,5]^{3}$.
Proof.
By Theorem 4.1, under (B2)–(B3), each confidence-weighted mean (i.e., $\textit{Novelty}_{i}$, $\textit{Contribution}_{i}$, $\textit{Feasibility}_{i}$) is an unbiased estimator of its true paper-level mean ($n_{i}^{*}$, $d_{i}^{*}$, $f_{i}^{*}$).
Assumption (B4) further ensures that these three means are mutually independent. Since the $p$-norm and normalization are deterministic and monotonic, the final innovation score $\hat{\textit{Innovation}}_{i}^{(p)}$ is thus an unbiased and monotonic transformation of the true means. ∎
Summary. HSPIM+ provides an unbiased estimator for composite innovation in expectation, under the independence assumptions.
5 Experiment
@@
Table 2: Statistics of peer review datasets.
@@
5.1 Datasets
We choose scientific peer review datasets from ACL-2017, CoNLL-2016, and ICLR-2017 in PeerRead [14], and the COLING-2020 dataset from NLPeer [15]. Table 2 shows that ICLR-2017 belongs to the ML/AI domain, while the others are from the NLP/CL domain. As describe in Step-4 of Section 3.2, these datasets provide numeric attribute scores of originality and soundness/correctness to form the ground-truth innovation score. Note that originality scores are mainly available in NLP-related conferences. For ICLR-2017, originality scores were manually annotated by the PeerRead authors and validated through the consistency re-annotation experiment. All labeled scores range from 1 to 5 (integers only).
5.2 Metrics
@@
Table 3: Performance comparison among zero-shot HSPIM, supervised deep learning models, and other zero-shot LLMs. The bold values represent the best results.
@@
5.3 Baselines
5.4 Hyperparameters and Experimental Setting
Genetic Algorithm Hyperparameters: To test the effectiveness under resource-limited scenarios, we set the number of iterations $I=5$ and the population size $P$ to 10. The predefined specific and common question sets each contain 11 candidates. The initial question-prompt combination used in naive HSPIM is selected from these sets. We set the mutation rate to 10%, and the number of elite individuals to $\max\{1,\,0.2\times P\}$. The fitness function uses RMSE to evaluate prediction performance.
Supervised Baseline Settings: We use the PeerRead setup for deep learning models and take the first 1000 tokens of each paper as input. We select only the originality and soundness/correctness attribute scores for training. The training epochs are set to 10 for CNN, LSTM, and Transformer. For CNN and LSTM, we set the number of layers to 2, the dropout rate to 0.5, the learning rate to 0.001, and the batch size to 64. For the Transformer encoder, we use 12 attention heads, a hidden size of 768, a learning rate of 2e-5, and a batch size of 8. All results are averaged over three runs.
Table 4: Performance comparison on ICLR-2017, CoNLL-2016, and ACL-2017: zero-shot HSPIM vs. DL models trained on COLING-2020. The bold values represent the best results.
@@ 5.4 Hyperparameters and Experimental Setting
5.5 Main Results
This section presents the main experimental results. We compare our HSPIM framework with baseline models on four datasets using RMSE and MAE as evaluation metrics. In the following results, HSPIMtwo-step and HSPIMjoint denote the two-step and joint optimization for HSPIM, respectively. HSPIMpruning=3 selects three section types ($s^{\prime}=3$) to compute the innovation scores, following Section 3.4. We use HSPIMjoint as the default framework to represent HSPIM with prompt optimization in the following experiments.
5.5.1 Zero-shot HSPIM vs. Supervised DL Models and Zero-shot LLMs
Table 3 shows that HSPIM framework outperforms both supervised baseline models and zero-shot baseline LLMs in terms of RMSE and MAE. Compared to one-step zero-shot prompting on full papers, HSPIM based on paper-to-sections-to-QAs decomposition reduces the average RMSE by 35.4% (from 0.916 to 0.592) for DeepSeek and by 17.3% (from 0.770 to 0.634) for GPT. Besides, our HSPIM outperforms all supervised baselines. This proves that the zero-shot approach can generalize well. For instance, DeepSeek-HSPIMpruning=3 achieves a 6.2% performance improvement over the supervised Transformer.
Additionally, we observe the following: (i) DeepSeek-HSPIM achieves better overall performance than GPT-HSPIM, which is reasonable because DeepSeek-V3 has stronger fundamental capabilities. (ii) HSPIMpruning=3 achieves the best results among the three optimization strategies. Fig. 6 shows the top-3 selected section types across different datasets and LLMs. We observe that Abstract, Analysis Theory, Experiment Analysis, and Conclusion frequently appear. It indicates these sections contain important novelty explanations and method contributions for measuring innovation. We believe section pruning helps to filter out potential noisy section types. In addition, it guides human reviewers to focus on specific sections that most significantly contribute to innovation analysis. (iii) HSPIM achieves relatively lower performance on the ACL-2017 dataset due to dataset differences. ACL-2017 has a higher average ground-truth innovation score (3.97) on the test set than ICLR-2017 (3.57), CoNLL-2016 (3.48), and COLING-2020 (3.44). Section 6.5 presents an analysis of the ground-truth innovation score distributions. Supervised baselines perform better since both training and test sets in ACL-2017 have high scores. LLM baselines (DeepSeek-V3 and GPT-4o mini) also perform well on ACL-2017 because LLMs tend to assign higher scores to longer texts [34]. This tendency matches the scoring distribution of ACL-2017. However, LLM baselines perform poorly on other datasets.
5.5.2 Zero-Shot HSPIM vs. DL Models (Transfer from COLING-2020)
Table 4 compares HSPIM with DL models trained on COLING-2020 and evaluated on the remaining three datasets. This setting provides a fairer comparison. HSPIM significantly outperforms DL models on ICLR-2017, CoNLL-2016, and ACL-2017. Compared to the best results of CNN, LSTM, and Transformer, HSPIM achieves average improvements of 44.0%, 41.8%, and 33.1%, respectively. These results confirm the model generalization ability of the zero-shot HSPIM framework.
Table 5: RMSE comparison across our proposed frameworks. The bold values represent the best results.